2m3 of soil containing 35% sand was mixed into 6m3 of soil containing 15% sand.What is the sand content of the mixture?
Answer by oberobic(2304)

2m^3 @ 35% sand = .35*2 = .7 m^3 of sand
6m^3 @ 15% sand = .15*6 = .9 m^3 of sand
.
So we have 8 m^3 of soil containing 1.6 m^3 of sand.
.
1.6/8 * 100 = 20% sand by volume
.
Done.